Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Difficult Math Problem #87 - Probability

Expert replies
by 800guy » Wed Jan 17, 2007 11:33 am
As a part of a game, four people each much secretly chose an integer between 1 and 4 inclusive. What is the approximate likelihood that all four people will chose different numbers?

this question comes from the difficult math doc on this website, OA coming after some people answer.
Join the discussion
Source: — Problem Solving |

by maxim730 » Wed Jan 17, 2007 3:14 pm
4!?
Join the discussion

by sumitkrishna » Thu Jan 18, 2007 9:08 pm
1/24 ???
Join the discussion

by sarika.k » Thu Jan 18, 2007 9:36 pm
3/32 ?
Join the discussion

by aim-wsc » Fri Jan 19, 2007 1:56 am
c'mon guys come up with explanation.

why do you think so?
Join the discussion

by thankont » Fri Jan 19, 2007 3:45 am
Choosing the same number is (1/4)^4 since say our chosen # is 2 then prob. of first person choosing 2 is 1/4 etc.
Now here we want different numbers so let first person choose 1 with prob. (1/4), second chooses number 2 with p = 1/4 so I think that probability is the same with the first case and p=(1/4)^4
Join the discussion

OA

by 800guy » Fri Jan 19, 2007 9:31 am
i second aim-wsc's comments..the purpose of sharing these questions is to get more discussion around explanations, sharing of knowledge.

here's the OA:

The probability that the first person will pick unique number is 1 (obviously) then the probability for the second is 3/4 since one number is already picked by the first, then similarly the probabilities for the 3rd and 4th are 1/2 and 1/4 respectively. Their product 3/4*1/2*1/4 = 3/32
Join the discussion

by mankey » Wed Sep 21, 2011 9:27 am
Favourable cases: 4! (All assigned different numbers)
Total cases: 4^4 (Since each number can be assigned to any person)

Probability=4!/4^4=3/32

Please let me know if this approach is correct?

Thanks
Mayank
Join the discussion

by GMATGuruNY » Wed Sep 21, 2011 10:23 am
800guy wrote:As a part of a game, four people each much secretly chose an integer between 1 and 4 inclusive. What is the approximate likelihood that all four people will chose different numbers?

this question comes from the difficult math doc on this website, OA coming after some people answer.
The first person can choose any of the four numbers.

P(2nd person chooses a number not yet chosen) = 3/4. (4 total numbers, 3 of them not yet chosen.)
P(3rd person chooses a number not yet chosen) = 2/4. (4 total numbers, 2 of them not yet chosen.)
P(4th person chooses a number not yet chosen) = 1/4. (4 total numbers, 1 of them not yet chosen.)

Since we want all of the events above to happen together, we multiply the fractions:
3/4 * 2/4 * 1/4 = 3/32.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by sl750 » Wed Sep 21, 2011 11:47 am
I am just curious. If the first number is no longer available for picking, the next person has to chose from the remaining 3 numbers, so how is that the total numbers still remain as 4 and not 3?
Join the discussion